OpenStudy (anonymous):

use algebra to solve each equation over the doman 0< x < 2pi give roots to nearest hundredth when necessary cos x + root 3 sinx = 0

OpenStudy (dan815):

u are preettyy u dont need math

OpenStudy (agent0smith):

\[\large \sqrt 3 sinx = -\cos x\] divide both sides by cosx so sinx becomes cosx \[\large \sqrt 3 \tan x = -1\] now divide both sides by sqrt3 and you'll have\[\large \tan x = \frac{ -1 }{\sqrt3 }\]

OpenStudy (agent0smith):

***sinx becomes tanx

OpenStudy (anonymous):

oh okay thank you! i remember it now! thanks

OpenStudy (dan815):